OpenStudy (ryan0716):

negative 1 over 2 x + 1 = –x + 8

OpenStudy (♪chibiterasu):

Please post this in the Mathematics section: http://openstudy.com/study#/groups/Mathematics You will receive quicker help by posting in the right section. :)

pooja195 (pooja195):

Hi! Welcome to openstudy :] You are indded in the wrong section but i'll help! :) Start by subtracting 1 from both sides we do this because we are isolating x \[\huge~\rm\frac{ 1 }{ 2 } x+1-1=-x+8-1\]

OpenStudy (♪chibiterasu):

Yeah, but the 1/2 is negative :p

pooja195 (pooja195):

;p oops

OpenStudy (anonymous):

i got it

OpenStudy (anonymous):

Rearrange the equation by subtracting what is to the right of the equal sign from both sides of the equation : -1/2*x+1-(-x+8)=0

OpenStudy (anonymous):

Step 1 : 1 Simplify — 2

OpenStudy (♪chibiterasu):

He's already receiving help: http://openstudy.com/study#/updates/56546ee8e4b0b09ed5dcfa6d

OpenStudy (anonymous):

step 2 Rewrite the whole as a fraction using 2 as the denominator : 1 1 • 2 1 = — = ————— 1 2 Combine the numerators together, put the sum or difference over the common denominator then reduce to lowest terms if possible: -x + 2 2 - x —————— = ————— 2 2 step 3 Rewrite the whole as a fraction using 2 as the denominator : 8 - x (8 - x) • 2 8 - x = ————— = ——————————— 1 2 Add up the two equivalent fractions (2-x) - ((8-x) • 2) x - 14 ——————————————————— = —————— 2 2 When a fraction equals zero ... Where a fraction equals zero, its numerator, the part which is above the fraction line, must equal zero. Now,to get rid of the denominator, Tiger multiplys both sides of the equation by the denominator. Here's how: x-14 ———— • 2 = 0 • 2 2 Now, on the left hand side, the 2 cancels out the denominator, while, on the right hand side, zero times anything is still zero. The equation now takes the shape : x-14 = 0 Solve : x-14 = 0 Add 14 to both sides of the equation : x = 14 the answer would be x = 14

OpenStudy (♪chibiterasu):

Who is tiger

OpenStudy (anonymous):

i use tiger algebra

OpenStudy (anonymous):

look it will explain it to you but i guess it is easier to put it in chat instead of typing it all.

OpenStudy (athecool):

Doesn't this belong in another section?